# Johannes Matzat        Aiko Schmeißer

### v4

@@ B.1 Union Election Data
Data sources. We start by accessing data on NLRB union representation elections between 1961 and 2009 from the replication package of Knepper (2020). The data were originally compiled by Farber (2016). Then, we add data on elections between 2010 and 2018 from NLRB election reports available on https://www.nlrb.gov/reports/agency-performance/election-reports. Together, our data cover the universe of union elections between 1961 and 2018 and includes information on vote counts, voting outcome, petition filing and election date, establishment name, address, and industry, as well as the name of the union organization.
@@
  • For multiple entries that reflect elections where more than one union was on the ballot or where different worker groups formed different bargaining units, we follow Frandsen (2021) and retain only the entry with the largest union vote share.
@@
B.2 Details on the Matching of Elections and Campaign Contributions
@@
Geocode commuting zones. In preparation for the spatial match, we first geocode all union election establishments based on their city and state (using the Open Street Map and Google Maps APIs) and assign the 1990 commuting zone. For the employees’ campaign contributions, we rely on donor addresses geocoded by Bonica (2019) up to 2016.484848Bonica (2019) contains campaign contributions until 2018 but geocodes are only provided until 2016. We use these geocodes to match to them the 1990 commuting zones.
@@
Firm name cleaning. Firm names in both the union election and the contribution data are cleaned and harmonized using the stnd_compname Stata command developed by Wasi and Flaaen (2015). The algorithm removes non-standard characters and whitespaces, doing-as-business and FKA names, as well as business entity types (e.g., CORP, INC, LLC). Moreover, it abbreviates common strings in firm names (e.g., Manufacturing $\rightarrow$ MFG, Professional $\rightarrow$ PROF).
@@
Linkage algorithm. For each commuting zone, we create lists of all cleaned firm names from the union election and the contribution data. Then, we use the reclink2 Stata command from Wasi and Flaaen (2015) to compare the string similarity of firm names.494949reclink2 builds on reclink written by Blasnik (2010). For each possible pair of firm names within the commuting zone, the command computes modified bigram scores. We keep potential matches with a score of at least .98 and manually review all of them. We identify roughly 70% of them as correct matches. [^50] The share of matches identified as correct is strongly increasing in the bigram score. For scores between .995 and 1, we keep 90% of the potential matches, while for scores between .98 and .985 this share is only 34%. We also tried keeping potential matches with a lower score (.95), but a manual review of a subsample of those revealed that a very low share of them represented correct matches. In our review, we generally took a conservative approach and were more tolerant of possibly rejecting a true match than retaining an incorrect match. This means that we measure a lower bound for the sum of contributions from all employees of an establishment. To demonstrate the spatial dimension of the matching procedure, Figure B.1 shows an example for the location of a union election establishment and all campaign contributions matched to it.
@@
Establishment-level aggregation. As a last step, we use all contributions with a matched establishment name and sum them up at the establishment-election cycle level. Our period of analysis covers three cycles before to three cycles after each union election, i.e., we observe each establishment over a period of seven cycles (14 years). While we generally keep establishment-cycle observations without any matched contribution and code them as zero, we retain only establishments for which we observe at least one matched contribution over the 14-year period. Out of the initial 28,823 union election establishments, we thereby keep 6,063 matched establishments which form our final estimation sample. Table A.1 compares the characteristics of matched and non-matched establishments.
@@ B.3 Occupation Classification
Occupation harmonization. The free-text occupations reported in DIME are not standardized. Thus, we map them to the 6-digit Standard Occupation Classification. For this, we combine an ensemble classifier called SOCcer (Russ et al., 2016), fuzzy string matching to an extensive crosswalk of laymen’s occupation titles from O*NET, as well as manual reviews from Dreher et al. (2023) and manual reviews of the most common occupation titles. In particular, we implement the following steps to identify good matches between a free-text occupation and a SOC code. First, we keep a match determined by SOCcer if the score of the first best match is higher than 0.3 and the difference to the second best match is larger than 0.1. Secondly, we search for exact matches of any substring of the free-text occupations and a list of laymen’s occupation titles, abbreviations and reported titles by experts obtained from O*NET. Thirdly, we fuzzy match the lists from O*NET with the free-text occupations and keep matches with a score above 0.99. Fourthly, we add matches from Dreher et al. (2023), which are based on a manual review. Finally, we manually review the free-text occupations that appear more than 50 times in our database of candidate contributions. With that procedure, we are able to assign a SOC code to 72% of all candidate contributions in our matched sample.
@@
Appendix C Effects of Losing a Union Election
@@
We estimate the effects of losing a union election compared to holding no election by using establishments who hold and lose an election in the future as a control group. Consider the treatment cohort of elections that were held and lost in the cycle 1985/86. Given that we observe each establishment only up to three cycles before the union election, we can use elections held and lost in the next two cycles as control cohorts. The untreated pre-election observations of the 1987/88 control cohort refer to the cycles 1981/82, 1983/84, and 1985/86 (event times $k=\{-2,-1,0\}$ of the treated cohort), and those of the 1989/90 control cohort refer to the cycles 1983/84, 1985/86, and 1987/1988 (event times $k=\{-1,0,1\}$ of the treated cohort). Note that later cohorts are not observed before the treated cohort hold their election and can therefore not be used in a DiD comparison. Consequently, we only have untreated observations that we can compare to the treated cohort’s observations in cycles 1981/82, 1983/84, 1985/86, and 1987/88 (event times $k=\{-2,-1,0,1\}$). This means we can only identify short-term effects.
@@
where $k$ now denotes the number of cycles relative to the cycle when the treated cohort held its union election. Establishment fixed effects are now saturated with indicators for the cohort-specific dataset $c$ to account for the fact that establishments enter several datasets. The DiD coefficient $\delta_{\textrm{DiD}}$ is given by the interaction between a dummy for post-election cycles of the treated cohort ($k\geq 0$) and a dummy for the treated cohort ($c_{i}=c$). Results are reported in Panel A of Table A.9.
In Panels B and C of Table A.9, we also show results for the alternative staggered DiD estimators by Borusyak et al. (forthcoming) and Callaway and Sant’Anna (2021). In line with our stacking implementation, in settings with no never-treated units, both estimators use not-yet-treated observations as controls. The methods differ from the stacked DiD model in the number of pre-treatment periods used and the aggregation of unit- or cohort-specific effects. In our results, however, the estimates are very similar to those of the stacked DiD model.

### v5

@@ B.1 Union Election Data
Data sources. We start by accessing data on NLRB union representation elections between 1961 and 2009 from the replication package of Knepper2020. The data were originally compiled by Farber2016. Then, we add data on elections between 2010 and 2018 from NLRB election reports available here. Together, our data cover the universe of union elections between 1961 and 2018 and includes information on vote counts, voting outcome, petition filing and election date, establishment name, address, and industry, as well as the name of the union organization.
@@
  • For multiple entries that reflect elections where more than one union was on the ballot or where different worker groups formed different bargaining units, we follow Frandsen2020 and retain only the entry with the largest union vote share.
@@
B.2 Matching of Union Elections and Campaign Contributions
@@
Geocode commuting zones. In preparation for the spatial match, we first geocode all union election establishments based on their city and state (using the Open Street Map and Google Maps APIs) and assign 1990 commuting zones. For the employees’ campaign contributions, we rely on donor addresses geocoded by Bonica2019 up to 2016.494949Bonica2019 contains campaign contributions until 2018 but geocodes are only provided until 2016. We use these geocodes to match to them 1990 commuting zones.
@@
Firm name cleaning. Firm names in both the union election and the contribution data are cleaned and harmonized using the stnd_compname Stata command developed by Wasi2015. The algorithm removes non-standard characters and whitespaces, doing-as-business and FKA names, as well as business entity types (e.g., CORP, INC, LLC). Moreover, it abbreviates common strings in firm names (e.g., Manufacturing $\rightarrow$ MFG, Professional $\rightarrow$ PROF).
@@
Linkage algorithm. For each commuting zone, we create lists of all cleaned firm names from the union election and the contribution data. Then, we use the reclink2 Stata command from Wasi2015 to compare the string similarity of firm names.505050reclink2 builds on reclink written by Blasnik2010. For each possible pair of firm names within the commuting zone, the command computes modified bigram scores. We keep potential matches with a score of at least .98 and manually review all of them. We identify roughly 70% of them as correct matches. [^51] The share of matches identified as correct is strongly increasing in the bigram score. For scores between .995 and 1, we keep 90% of the potential matches, while for scores between .98 and .985 this share is only 34%. We also tried keeping potential matches with a lower score (.95), but a manual review of a subsample of those revealed that a very low share of them represented correct matches. In our review, we generally took a conservative approach and were more tolerant of possibly rejecting a true match than retaining an incorrect match. This means that we measure a lower bound for the sum of contributions from all employees of an establishment. To demonstrate the spatial dimension of the matching procedure, Figure B.1 shows an example for the location of a union election establishment and all campaign contributions matched to it.
GPT vs. manual linkage comparison. The rapid advancements in large language models (LLMs) have made alternative linkage procedures feasible that were unavailable when we initiated this project. Specifically, the low cost of going through a large number of potential matches possibly changes the cost-benefit calculation that is considered when setting the reclink2 threshold above which we review potential matches. To test the benefits of lowering the threshold and employing an LLM-based approach, we draw a random sample of 100 union elections, compute modified bigram scores again between all potential combinations of firm names in the union election and contribution data within the same commuting zone, and keep all potential matches with a score of at least 0.5. This yields 10,954 firm name pairs. We send these pairs to OpenAI’s API and instruct it to identify correct matches using the GPT-4o model. To establish the ground truth, we also manually review all 10,954 potential matches. Finally, we merge our original classification with the test sample for comparison. The LLM-based approach fails to improve the accuracy of our linkage procedure. The original procedure yields 99.2% agreement with the ground truth, while the LLM-based procedure yields 99.1%. [^52] To assess the quality of the linkage procedure, we also calculate Cohen’s Kappa (cohen1960coefficient), which takes on the value 0 when the agreement amount is expected by chance and 1 when there is perfect agreement. Both procedures achieve a Kappa score of 0.9, indicating an “almost perfect” level of agreement with the ground truth, as defined by landis1977one; landis1977measurement. Although the LLM classifies the vast majority of cases correctly, it misclassifies some matches as correct, which outweighs its benefit in detecting rare matches with a bigram score below 0.98.
@@
Establishment-level aggregation. As a last step, we use all matched contributions and sum them up at the establishment-election cycle level. Our period of analysis covers three cycles before to three cycles after each union election, i.e., we observe each establishment over a period of seven cycles (14 years). While we generally keep establishment-cycle observations without any matched contribution and code them as zero, we retain only establishments for which we observe at least one matched contribution over the 14-year period. Out of the initial 28,823 union election establishments, we thereby keep 6,063 matched establishments which form our final estimation sample. Table A.1 compares the characteristics of matched and non-matched establishments.
@@ B.3 Occupation Classification
Occupation harmonization. The free-text occupations reported in DIME are not standardized. Thus, we map them to the 6-digit Standard Occupation Classification. For this, we combine an ensemble classifier called SOCcer (Russ2016), fuzzy string matching to an extensive crosswalk of laymen’s occupation titles from O*NET, as well as manual reviews from Dreher2020 and manual reviews of the most common occupation titles. In particular, we implement the following steps to identify good matches between a free-text occupation and a SOC code. First, we keep a match determined by SOCcer if the score of the first best match is higher than 0.3 and the difference to the second best match is larger than 0.1. Secondly, we search for exact matches of any substring of the free-text occupations and a list of laymen’s occupation titles, abbreviations and reported titles by experts obtained from O*NET. Thirdly, we fuzzy match the lists from O*NET with the free-text occupations and keep matches with a score above 0.99. Fourthly, we add matches from Dreher2020, which are based on a manual review. Finally, we manually review the free-text occupations that appear more than 50 times in our database of candidate contributions. With that procedure, we are able to assign a SOC code to 72% of all candidate contributions in our matched sample.
@@
Alternative residence-based classification. Given the share of occupations that cannot be classified, as a robustness check we draw on an alternative classification that uses information on donors’ exact location of residence. Bonica2019 geocodes donor addresses in the DIME data, which we link to median household income at the census tract level using data from the 1980, 1990, and 2000 Censuses, as well as the 5-year American Community Survey samples from 2010 onward.. With that, we identify managers (workers) as donors who live in a census tract that has a median income above (below) the 80th or 90th percentile of the state-specific distribution of census-tract median incomes.
B.4 Measures of Labor Relations at the Establishment
Unfair labor practice charges. We collect data on unfair labor practice (ULP) charges reported to the NLRB from the CHIPS archive for 1984-2000 (compiled by Forest Gregg and available here), from the CATS archive for 1999-2011 (available here), and from the NLRB website for 2007-2020 (available here). The charges include complaints against employers for restraining employees in their rights to organize in a union or collectively bargain (NLRA § 8(a)(1)), dominating or controlling a union (§ 8(a)(2)), discharging or otherwise discriminating workers involved in organizing (§§ 8(a)(3) and 8(a)(4)), and failing to bargain in good faith with the union (§ 8(a)(5)). To focus on ULP committed by employers in the process of the organizing drive, we limit the data to charges related to violations of §§ 8(a)(1), 8(a)(3), and 8(a)(4).
The data contains the name, state, city, and address of the employer. Firm name and address are cleaned and harmonized using the Stata commands stnd_compname and stnd_address, respectively (Wasi2015). Then, we match a ULP charge to a union election in our estimation sample if one of the following conditions is fulfilled.
  • Exact match of state, firm name, and city
  • Exact match of state, firm name, and address
  • Exact match of state, city, and address
  • Exact match of state and match score above 0.9 from fuzzy merge on firm name, city, and address using reclink2 (if address information is missing in both datasets or not missing in both datasets)
  • Exact match of state and match score above 0.8 from fuzzy merge on firm name, city, and address using reclink2 (if address information is missing in one of both datasets)
Since charges must be filed within six months of an alleged violation, we identify all ULP charges that were filed between six months before to six months after a union election. With that, we find that 44% of all union elections in our sample involve a ULP charge, similar to the shares reported by Bronfenbrenner2009 and McNicholas2019.
Collective bargaining contracts. We combine bargaining contract data compiled by Holmes2006 for 1985-2003 (available here) and by Gregg2024 for 2004-2020 (available here). The data contains notices about both initial contracts, i.e. first-time negotiations after an election, and about the renegotiation or reopening of existing contracts. It is reported to the Federal Mediation and Conciliation Service (FMCS) for it to be able to prepare mediation services for the negotiation of contracts.
In the data, we observe the name, state, city, and street of the reporting establishment. We clean this information and match contract notices to establishments in our union election estimation sample using the same algorithm as above for the matching of ULP charges. We find a contract expiration notice in the five years following the union election for 48% of the won union elections in our sample, which is comparable to the numbers presented in DiNardo2004, Ferguson2008, and Frandsen2020.
Appendix C Instrumental Variable Strategy
In this section, we describe our IV approach, which complements the DiD strategy with arguably exogenous variation in union support driven by spikes in work-related fatalities. Specifically, we use as instrument sector-level fatal work accidents in the 30 days before a union election. [^56] The median time between petition and union election in our sample is 45 days. Less than 5% of all elections are held within 30 days after the petition. Thus, for almost all elections, the election date is fixed during our time of instrument exposure. Safety at work is a fundamental concern to all workers, especially when one’s life is in danger. Work-related fatalities are still common in the United States. In 2023, the Occupational Safety and Health Administration (OSHA) reported 5,283 deaths at work, more than 14 per day on average. Workplace safety is a central theme in union campaigns, and unions can be successful in improving safety conditions (e.g., AFLCIO2022; Hagedorn2016; Li2022). It is therefore plausible that increased attention to serious workplace hazards – such as fatal accidents – leads to a surge in support for unionization.
We implement the IV approach by estimating the following two-stage model:
where $\Delta y_{j}$ denotes the change in campaign contribution patterns in the three cycles after the union election relative to the three cycles before (excluding the cycle of the union election). By using changes as the outcome variable, the specification builds on the DiD approach and accounts for time-invariant differences between establishments that may affect the level of campaign contributions. Our main instrument is $A_{st}$, which represents the number of fatal accidents in 2-digit sector $s$ in the 30 days prior to the election after accounting for seasonal variation. [^57] Data on fatal work accidents is obtained from OSHA in the form of Fatality and Catastrophe Investigation Summaries (OSHA form 170). Figure C.1 depicts the exploited time variation. We allow for a non-linear effect by including $A_{st}^{2}$ and for a larger impact of fatalities in sectors where fatalities are common and where workers may be more concerned about workplace safety by the interaction term $A_{st}\times FR_{s}$ ($FR_{s}$ denotes the share of fatal work accidents occurring in a given sector out of all fatal work accidents in the sample). Importantly, instead of directly instrumenting union victory in a standard 2SLS approach, the first stage explains the continuous pro-union vote share $V_{j}$. In the second stage, we then use an indicator for predicted victory that is based on the predicted vote share in the first stage. This approach resembles the treatment assignment process and exploits the maximum available information. To account for the uncertainty from the first-stage regression, we compute standard errors by bootstrapping the two-stage model. The model includes several control variables: the main effect of $FR_{s}$, the number of fatalities and employees at the sector-year level (to account for sector-specific fatality trends), the number of eligible voters in the union election (as precision control), as well as year fixed effects $\gamma_{t}$ and month-of-the-year fixed effects $\mu_{m}$.
Results are reported in Table C.1. The first-stage results show that sector-level fatal work accidents are a significant predictor of the union election outcome, with an F-statistic of 16.5. We find that the positive effect of spikes in work accidents on unionization is stronger in sectors where work accidents are more common, i.e., where workplace safety may be a greater concern for workers. The second-stage results confirm our main findings from the DiD model, highlighting a leftward shift in campaign contributions in response to unionization. The magnitude of the coefficients is comparable but slightly larger than in the DiD model. As compliers respond to information on fatal work accidents, we deem it plausible that they also react more strongly to information provided by unions and to changes to their work environment induced by unionization. The estimates are considerably less precise, however. While the effects on the party composition of contributions from managers are still significant at the 5% level, the effects for workers are no longer significant. [^58] We also verify the IV approach with a falsification exercise. We re-estimate model (9) using the change in campaign contribution patterns between t1t-1 and t2t-2 as the outcome. We do not find any evidence for pre-existing differential trends related to spikes in fatal work accidents.
Figure C.1: Seasonally Adjusted Fatal Work Accidents, 1984-2012
Table C.1: Instrumental Variable Results
IHS(\$ to all candidates) IHS(\$ to Dem.) - IHS(\$ to Rep.)
All Workers Managers All Workers Managers
(1) (2) (3) (4) (5) (6)
[A] OLS
$\mathbbm{1}[V_{j}>.5]$ -0.092 0.038 -0.072 0.227∗∗∗ 0.089∗∗ 0.232∗∗∗
(0.082) (0.044) (0.062) (0.079) (0.041) (0.056)
[B] 2nd stage
$\mathbbm{1}[\widehat{V_{j}}>.5]$ 0.036 0.086 -0.042 0.334 0.115 0.260∗∗
(0.174) (0.097) (0.134) (0.176) (0.086) (0.125)
[C] 1st stage
$A_{st}$ 0.002 0.002 0.002 0.002 0.002 0.002
(0.007) (0.007) (0.007) (0.007) (0.007) (0.007)
$A_{st}^{2}$ -0.003∗∗ -0.003∗∗ -0.003∗∗ -0.003∗∗ -0.003∗∗ -0.003∗∗
(0.001) (0.001) (0.001) (0.001) (0.001) (0.001)
$A_{st}\times FR_{s}$ 0.223∗∗∗ 0.223∗∗∗ 0.223∗∗∗ 0.223∗∗∗ 0.223∗∗∗ 0.223∗∗∗
(0.055) (0.055) (0.055) (0.055) (0.055) (0.055)
K-P F-stat 16.50 16.50 16.50 16.50 16.50 16.50
[D] 2nd stage falsification: pre-trend
$\mathbbm{1}[\widehat{V_{j}}>.5]$ -0.007 0.093 0.033 0.124 -0.020 0.046
(0.207) (0.094) (0.116) (0.230) (0.100) (0.129)
Notes: The table reports IV results for the effect of unionization on the IHS-transformed total amount contributed (columns (1) - (3)) and on the difference between the IHS-transformed amounts contributed to Democratic and Republican candidates (columns (4) - (6)). Panel A reports OLS coefficients, Panel B reports the second-stage coefficients from model (9), and Panel C reports the first-stage coefficients from model (8). In Panels A and B, the outcome is the difference between the average outcome in the three cycles after and the average outcome in the three cycles before the union election (excluding the cycle of the union election). In Panel D, the outcome is the change between one and two cycles before the union election. $N=5,803$ establishments. Bootstrapped standard errors (with 500 replications), shown in parentheses, are clustered at the establishment level. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Appendix D Effects of Losing a Union Election
@@
We estimate the effects of losing a union election compared to holding no election by using establishments that hold and lose an election in the future as a control group. Consider the treatment cohort of elections that were held and lost in the cycle 1985/86. Given that we observe each establishment only up to three cycles before the union election, we can use elections held and lost in the next two cycles as control cohorts. The untreated pre-election observations of the 1987/88 control cohort refer to the cycles 1981/82, 1983/84, and 1985/86 (event times $k=\{-2,-1,0\}$ of the treated cohort), and those of the 1989/90 control cohort refer to the cycles 1983/84, 1985/86, and 1987/1988 (event times $k=\{-1,0,1\}$ of the treated cohort). Note that later cohorts are not observed before the treated cohort hold their election and can therefore not be used in a DiD comparison. Consequently, we only have untreated observations that we can compare to the treated cohort’s observations in cycles 1981/82, 1983/84, 1985/86, and 1987/88 (event times $k=\{-2,-1,0,1\}$). This means we can only identify short-term effects.
@@
where $k$ now denotes the number of cycles relative to the cycle when the treated cohort held its union election. Establishment fixed effects are now saturated with indicators for the cohort-specific dataset $c$ to account for the fact that establishments enter several datasets. The DiD coefficient $\delta_{\textrm{DiD}}$ is given by the interaction between a dummy for post-election cycles of the treated cohort ($k\geq 0$) and a dummy for the treated cohort ($c_{i}=c$). Results are reported in Panel A of Table D.1.
In Panels B and C of Table D.1, we also show results for the alternative staggered DiD estimators by Borusyak2021 and Callaway2021. In line with our stacking implementation, in settings with no never-treated units, both estimators use not-yet-treated observations as controls. The methods differ from the stacked DiD model in the number of pre-treatment periods used and the aggregation of unit- or cohort-specific effects. In our results, however, the estimates are very similar to those of the stacked DiD model.
Table D.1: Effects of Losing a Union Election
IHS(\$ to all candidates) IHS(\$ to Dem.) - IHS(\$ to Rep).
All Workers Managers All Workers Managers
(1) (2) (3) (4) (5) (6)
[A] Stacking
$\delta_{\textrm{DiD}}$ -0.049 -0.026 0.070 0.057 -0.013 0.037
(0.088) (0.040) (0.053) (0.097) (0.043) (0.057)
N 31,501 31,501 31,501 31,501 31,501 31,501
[B] Borusyak, Jaravel, and Spiess (2021)
$\delta_{\textrm{DiD}}$ -0.048 -0.029 0.075 0.080 -0.007 0.048
(0.090) (0.045) (0.059) (0.100) (0.049) (0.064)
N 16,658 16,658 16,658 16,658 16,658 16,658
[C] Callaway and Sant’Anna (2021)
$\delta_{\textrm{DiD}}$ -0.043 -0.038 0.061 0.076 -0.007 0.053
(0.095) (0.047) (0.064) (0.105) (0.052) (0.070)
N 16,658 16,658 16,658 16,658 16,658 16,658
Notes: The table presents DiD estimates for the effect of losing a union election versus holding no election. We compare establishments with a lost union election in a given cycle (treated cohort) with establishments with a lost union election in one of the next two cycles (control cohorts) in a DiD design. Thereby, we estimate short-term effects of losing an election (i.e., for event times $k=\{0,1\}$). Panel A shows results from a stacked DiD model, and Panels B and C implement the staggered DiD estimators of Borusyak2021 and Callaway2021. See Appendix D for details of the implementation. Standard errors clustered at the establishment level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$